# Seeing, Listening, Remembering, and Reasoning: A Multimodal Agent with Long-Term Memory

### v2

@@ Abstract
We introduce M3-Agent, a novel multimodal agent framework equipped with long-term memory. Like humans, M3-Agent can process real-time visual and auditory inputs to build and update its long-term memory. Beyond episodic memory, it also develops semantic memory, enabling it to accumulate world knowledge over time. Its memory is organized in an entity-centric, multimodal format, allowing deeper and more consistent understanding of the environment. Given an instruction, M3-Agent autonomously performs multi-turn, iterative reasoning and retrieves relevant information from memory to accomplish the task. To evaluate memory effectiveness and memory-based reasoning in multimodal agents, we develop M3-Bench, a new long-video question answering benchmark. M3-Bench comprises 100 newly recorded real-world videos captured from a robot’s perspective (M3-Bench-robot) and 920 web-sourced videos across diverse scenarios (M3-Bench-web). We annotate question-answer pairs designed to test key capabilities essential for agent applications, such as human understanding, general knowledge extraction, and cross-modal reasoning. Experimental results show that M3-Agent, trained via reinforcement learning, outperforms the strongest baseline, a prompting agent using Gemini-1.5-pro and GPT-4o, achieving 6.7%, 7.7%, and 5.3% higher accuracy on M3-Bench-robot, M3-Bench-web and VideoMME-long, respectively. Our work advances the multimodal agents toward more human-like long-term memory and provides insights into their practical design. Model, code and data are available at https://github.com/bytedance-seed/m3-agent.
@@ 1 Introduction
Imagine that in the future a household robot can autonomously carry out household tasks without your explicit instructions; it must have learned the operational rules of your home through daily experiences. In the morning, it hands you a cup of coffee without asking "coffee or tea?", because it has gradually formed a memory of you, tracking your preferences and routines through long-term interaction. For a multimodal agent, achieving such a level of intelligence fundamentally relies on three capabilities: (1) continuously perceiving the world through multimodal sensors; (2) storing its experiences in long-term memory and gradually building knowledge about the environment; (3) reasoning over this accumulated memory to guide its actions.
@@
  • Episodic memory: Records concrete events observed within the video. For example, "Alice takes the coffee and says, ‘I can’t go without this in the morning,’" and "Alice throws an empty bottle into the green garbage bin."
  • Semantic memory: Derives general knowledge from the clip. For example, "Alice prefers to drink coffee in the morning" and "The green garbage bin is used for recycling."
The generated memories are then stored in long-term memory, which supports multimodal information such as faces, voices and textual knowledge. Moreover, the memory is organized in an entity-centric structure. For example, information related to the same person (e.g., their face, voice and associated knowledge) is connected in a graph format, as shown in Figure 1. These connections are incrementally established as the agent extracts and integrates semantic memory.
@@
We evaluate M3-Agent on long video question answering (LVQA), where the videos simulate the multimodal input streams (visual and auditory) received by an agent. Most existing LVQA benchmarks [9, 60, 2, 48] mainly focus on visual understanding, such as action recognition and spatial/temporal perception, leaving a gap in evaluating higher-level cognitive abilities that rely on long-term memory and are crucial for real-world agents, such as understanding humans, extracting general knowledge, and performing cross-modal reasoning. To bridge this gap, we introduce M3-Bench, a new LVQA benchmark designed to evaluate a multimodal agent’s ability to reason with long-term memory. M3-Bench consists videos from two sources: (1) M3-Bench-robot, consisting of 100 real-world videos recorded from a robot’s perspective, and (2) M3-Bench-web, comprising 920 YouTube videos spanning a broader range of content and scenarios. We define five question types, as shown in Table 1, targeting different aspects of memory-based reasoning. In total, we annotate 1,276 QA pairs for M3-Bench-robot and 3,214 QA pairs for M3-Bench-web.
@@ 2.1 Long-Term Memory of AI Agents
However, most existing approaches focus on LLM agents. In contrast, multimodal agents process a broader range of inputs and store richer, multimodal content and concepts in memory. This also introduces new challenges, particularly in maintaining consistency of long-term memory. Moreover, just as humans acquire world knowledge through experience, multimodal agents should form internal world knowledge in memory, rather than merely storing description of experience.
@@ 2.2 Online Video Understanding
Multi-detail Reasoning This requires aggregating multiple pieces of information distributed across the video.
@@
Human Under- standing This involves reasoning about human-related attributes such as identity, emotions, personality, or relationships.
@@ 3 Datasets
#Videos Len.(s) #QAs Anno. Form. Agent Cross- Human Knowledge
@@
Table 2: Comparison of M3-Bench with existing long-video question answering benchmarks across key dimensions: number of videos (#Videos), average video length in seconds (Len.), number of QA pairs (#QAs), annotation method (Anno., M/A denote manually/automatic), question format (Form., O/C indicate open-ended/close-ended), presence of an agent in the video (Agent Present), inclusion of cross-modal reasoning questions (Cross-Modal QA), human understanding questions (Human QA), and questions about general knowledge (Knowledge QA).
@@ 3.1 M3-Bench-robot
Video Filming Recording videos with actual robots poses significant challenges due to high operational costs, hardware limitations, and deployment complexities. To address these constraints, we adopted a practical alternative: employing human actors to simulate robot behavior. This approach simplifies data collection while preserving both the first-person robot perspective and the multimodal quality required for our benchmark.
Each script involves multiple actors, with one designated to simulate the robot. This actor wore head-mounted camera equipment to capture the robot’s egocentric visual and auditory perspective. The resulting footage constituted the final videos in M3-Bench-robot. To ensure diversity and minimize location bias, we recruited 67 actors and filmed across 51 distinct locations, with no more than three videos recorded at each location.
We collected two types of audio tracks for each video. The first was directly recorded by the head-mounted device, reflecting the raw auditory input a robot would naturally receive, including ambient sounds and spatial acoustic variations. The second was captured using individual lapel microphones worn by each actor, providing high-fidelity voice recordings to complement the primary audio stream.
Annotations After recording the videos, annotators curated QA pairs for each video. Although some questions was pre-scripted, the final video content may deviate from the original script due to realistic filming conditions. Consequently, not all scripted questions remain applicable. Annotators carefully reviewed each scripted question to determine whether it should be retained, revised, or discarded, and provided corresponding answers when necessary. For all retained or revised questions, annotators were required to specify the precise timestamp at which the question should be asked. Importantly, the timestamp must precede the robot’s corresponding response or action to avoid inadvertently revealing the answer.
In addition to the script-based questions, annotators were also required to create new questions to ensure that each video contained at least 12 QA pairs. All newly added questions should also align with one or more of the question types listed in Table 1.
Besides QA pair creation, annotators also generated subtitles to enhance the usability of the dataset. Specifically, they manually annotated the start and end timestamps for each dialogue segment, together with the speaker’s identity and the transcribed dialogue content.
@@ 3.2 M3-Bench-web
Video Collection The video collection adopts a question-driven approach: annotators selected videos that could support the design of at least five questions belong to the types listed in Table 1. This strategy naturally led to the selection of videos with rich narratives and complex inter-entity relationships, making them well-suited for assessing agent’s capability of reasoning with long-term memory.
To promote video diversity and avoid overrepresentation of easily annotated content, we provided annotators with a reference list of video categories emphasizing high information density and relevance to real-world multimodal agent applications. Annotators are required to submit up to 20 videos from each category and are allowed to suggest new categories, which are included if deemed sufficiently distinct from the existing category list by the authors. The final dataset comprises 46 distinct video types, as summarized in Figure 3.
QA Annotations The same annotator who collected the video also generates at least five corresponding question-answer pairs. Each question must correspond to at least one type defined in Table 1. In M3-Bench-web, all question timestamps are set to the end of the video. All questions are required to be specific, objective, and have a single unambiguous answer that can be reasonably derived from clues in the video, ensuring both the effectiveness and fairness of subsequent evaluation. For example, questions answerable from multiple perspectives or with ambiguous references, such as "the man" or "in the middle part of the video," are not considered valid. Appendix 9 provides the full annotation guidelines, annotators’ information, and quality control details for M3-Bench-web.
@@ 3.3 Automatic Evaluation
We use GPT-4o as an automatic evaluator for M3-Bench by prompting it to assess the correctness of a generated answer by comparing it to the corresponding reference answer for the same question. The prompt template is shown in Table LABEL:prompt_gpt4o_evaluation (§ 15.1).
@@ 4 Approach
search_clip Retrieves memory from the top-kk relevant video clips for a text query, including episodic and semantic memory.
@@ 4.1 Long-Term Memory
Long-term memory is implemented as an external database that stores information in a structured, multimodal format (text, images, audio). Specifically, memory entries are organized as a memory graph, where each node represents a distinct memory item. Each node includes a unique ID, modality type, raw content, weight, embeddings, and other metadata such as timestamps. See Table 3 for details. Nodes are connected by undirected edges that capture logical relationships between memory items. These connections serve as threads that facilitate the retrieval of relevant memories.
The agent constructs its memory by incrementally adding new text, image, or audio nodes, along with edges connecting them, or by updating the content and weights of existing nodes. Conflicting information may be introduced during construction. To resolve this, M3-Agent applies a weight-based voting mechanism during inference: entries that are reinforced more frequently accumulate higher weights, allowing them to override conflicting entries with weaker reinforcement. This mechanism ensures the robustness and consistency of the memory graph over time.
Search Tool To facilitate memory access, we provide a suite of search tools that enable the agent to retrieve relevant memories based on specific requirements. In particular, we implement two types of search mechanisms operating at different levels of granularity, as summarized in Table 4. Detailed implementation of these retrieval mechanisms is provided in Appendix 10.
@@ 4.3 Control
When an instruction is received, the control process is triggered. As illustrated in Figure 1, during control, M3-Agent autonomously performs multi-turn reasoning and invokes search functions to retrieve relevant information from long-term memory, up to a maximum of HH rounds. M3-Agent can independently decide which search function to invoke, such as search_clip to retrieve specific memory clips, or search_node to obtain the character ID of a particular character.
@@ 4.4 Training
We apply reinforcement learning to optimize the M3-Agent. Although the memorization and control are conceptually handled by a single model, we trained two separate models to achieve optimal performance. Memorization relies strong multimodal understanding, while control requires strong reasoning capabilities. Accordingly, we initialized each model with different foundation models: Qwen2.5-Omni [49], an advanced open-source multimodal model supporting both visual and audio inputs, for memorization; and Qwen3 [51], an open-source large language model with powerful reasoning abilities, for control.
@@
Memorization To improve the model’s ability to generate desired memory, we perform imitation learning on Qwen2.5-Omni-7b to create memory-7b-sft. The process begins with constructing a high-quality synthetic demonstration dataset. We segment each video in the dataset into 30-second clips, and corresponding memory annotations are generated through a three-stage process: (1) Episodic memory synthesis: We perform a hybrid annotation strategy by jointly prompting Gemini-1.5-Pro and GPT-4o. Accordingly, GPT-4o supplies frame-level cues, which serve as priors for Gemini-1.5-Pro; the two outputs are merged to form richer narrative summaries than either alone. (2) Identity equivalence detection: We propose an algorithm that automatically mines high-confidence meta-clips, short monologue clips containing exactly one face and one voice, from a long video to construct a global face-voice correspondence. These meta-clips offer clear identity cues, enabling accurate face-voice pairing. Once the global mapping is established, it can be used to automatically annotate face-voice associations in any 30-second subclip. (3) Other semantic memory synthesis: We design prompt templates to extract semantic memories from various perspectives, guiding semantic memories to include information listed in Table 10 (§ 11). Details of the data synthesis process are provided in Appendix 11. In total, we synthesize 10,952 samples: 10,752 for training and 200 for validation.
@@ 5.1 Baselines
We evaluate MM-Agent against three types of baselines:
@@
  • Gemini-Agent: Gemini-1.5-Pro is prompted separately for memory access and control process. During memory access, it receives the full video with audio, facial recognition results and speaker identification results to generate episodic and semantic memories, denoted as memory-gemini-prompt. In the control process, it performs memory searches and generates responses, referred to as control-gemini-prompt.
  • Gemini-GPT4o-Hybrid: We also evaluate a setup where GPT-4o is prompted to perform memory search and generate responses (control-gpt4o-prompt). The memory access remains handled by memory-gemini-prompt.
@@ 5.3 Main Results
Table 5: Results on M3-Bench-robot, M3-Bench-web, and VideoMME-long. We also present a comparison of all methods across different question types in M3-Bench: multi-detail reasoning (MD), multi-hop reasoning (MH), cross-modal reasoning (CM), human understanding (HU), and general knowledge extraction (GK).
@@ 5.4 Ablation Study
To evaluate the impact of memorization on overall performance, we fixed the control model to control-7b-rl and compared different memorization methods, as shown in Table 6. First, we replaced the memory with that generated by memory-gemini-prompt, resulting in accuracy drops of 2.0%, 2.6%, and 9.1% on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. This suggests that memory-7b-sft produces higher-quality memory than memory-gemini-7b. Next, we evaluated memory-7b-prompt, which led to accuracy reductions of 5.4%, 9.0%, and 11.0% on the same benchmarks, highlighting the importance of imitation learning in generating effective memory. Finally, we ablated key components in the memory generation process. The results show that removing character identity equivalence or semantic memory significantly degrades QA performance.
@@
Next, we investigate the impact of the control on final performance. We fix memorization model as memory-7b-sft and evaluate various control process models, as shown in Table 7. First, we compare two RL algorithms: GRPO and DAPO. Training details for GRPO are provided in Appendix 13. Our results show that control-32b-rl trained with DAPO consistently outperform control-32b-grpo across all test sets. Second, we analyze how DAPO’s performance scales with model size. The results indicate substantial improvements across all sizes. Specifically, after DAPO training, control-32b-rl achieves improvements of 10.0%, 8.0%, and 9.3% in accuracy over control-32b-prompt on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. Finally, we ablate two designs: inter-instruction and reasoning. Both are shown to be critical. Removing inter-instruction results in accuracy drops of 10.5%, 5.8%, and 5.9% on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. Removing reasoning leads to decreases of 11.7%, 8.8%, and 9.5% on the same benchmarks.

### v4

@@ Abstract
We introduce M3-Agent, a novel multimodal agent framework equipped with long-term memory. Like humans, M3-Agent can process real-time visual and auditory inputs to build and update episodic and semantic memories, gradually accumulating world knowledge. Its memory is organized in an entity-centric, multimodal manner, enabling deeper and more consistent understanding of the environment. Given an instruction, M3-Agent autonomously performs multi-turn reasoning and retrieves relevant memories to complete tasks. To evaluate memory effectiveness and memory-based reasoning in multimodal agents, we develop M3-Bench, a long-video question answering benchmark comprising 100 newly recorded robot-perspective videos (M3-Bench-robot) and 920 diverse web-sourced videos (M3-Bench-web). We annotate QA pairs designed to test capabilities essential for agent applications, such as person understanding, general knowledge extraction, and cross-modal reasoning. Experimental results show that M3-Agent, trained via reinforcement learning, outperforms the strongest baseline, a prompting agent using Gemini-1.5-pro and GPT-4o, achieving 6.7%, 7.7%, and 5.3% higher accuracy on M3-Bench-robot, M3-Bench-web and VideoMME-long, respectively. Our work advances multimodal agents toward more human-like long-term memory and provides insights for their practical design. Model, code and data are available at https://github.com/bytedance-seed/m3-agent.
@@ 1 Introduction
Imagine that in the future a household robot can autonomously carry out household tasks without your explicit instructions; it must have learned the operational rules of your home through daily experiences. In the morning, it hands you a cup of coffee without asking “coffee or tea?", because it has gradually formed a memory of you, tracking your preferences and routines through long-term interactions. For a multimodal agent, achieving this level of intelligence fundamentally relies on three capabilities: (1) continuously perceiving the world via multimodal sensors; (2) storing and organizing its experiences into a long-term memory, and gradually building knowledge of its environment; (3) reasoning over this accumulated memory to guide its actions.
@@
  • Episodic memory: Concrete events observed within the video. For example, "Alice takes the coffee and says, ‘I can’t go without this in the morning,’" and "Alice throws an empty bottle into the green garbage bin."
  • Semantic memory: General knowledge from the clip. For example, "Alice prefers to drink coffee in the morning" and "The green garbage bin is used for recycling."
The generated contents are then integrated into the agent’s long-term memory, which supports multimodal information such as faces, voices, and textual knowledge. The memory is organized in an entity-centric structure. For example, information related to the same person (e.g., their face, voice, and associated knowledge) is connected within a graph, as shown in Figure 1. These connections are incrementally established as the agent extracts and integrates episodic and semantic memory.
@@
We evaluate M3-Agent on long video question answering (LVQA), where the videos simulate the multimodal input streams (visual and auditory) received by an agent. Most existing LVQA benchmarks [11, 62, 2, 50] mainly focus on visual understanding, such as action recognition and spatial/temporal perception, leaving a gap in evaluating higher-level cognitive abilities that rely on long-term memory and are crucial for real-world agents, such as understanding persons, extracting general knowledge, and performing cross-modal reasoning. To bridge this gap, we introduce M3-Bench, a new LVQA benchmark designed to evaluate a multimodal agent’s ability to reason with long-term memory. M3-Bench consists videos from two sources: (1) M3-Bench-robot, consisting of 100 real-world videos recorded from a robot’s perspective, and (2) M3-Bench-web, comprising 920 YouTube videos spanning a broader range of content and scenarios. We define five question types, as shown in Table 1, targeting different aspects of memory-based reasoning. In total, we annotate 1,276 QA pairs for M3-Bench-robot and 3,214 QA pairs for M3-Bench-web.
@@ 2.1 Long-Term Memory of AI Agents
However, most existing approaches focus on LLM agents. In contrast, multimodal agents process a broader range of inputs and store richer, multimodal content and concepts in memory [8, 7]. This also introduces new challenges, particularly in maintaining consistency of long-term memory. Moreover, just as humans acquire world knowledge through experience, multimodal agents should form internal world knowledge in memory, rather than merely storing description of experience.
@@ 2.2 Online Video Understanding
Multi-evidence Reasoning This requires aggregating multiple pieces of information distributed across the video.
@@
Person Under- standing This involves reasoning about person-related attributes such as identity, emotions, personality, or relationships.
@@ 3 Datasets
#Videos Len.(s) #QAs Anno. Form. Agent Cross- Person Knowledge
@@
Table 2: Comparison of M3-Bench with existing long-video question answering benchmarks across key dimensions: number of videos (#Videos), average video length in seconds (Len.), number of QA pairs (#QAs), annotation method (Anno., M/A denote manually/automatic), question format (Form., O/C indicate open-ended/close-ended), presence of an agent in the video (Agent Present), inclusion of cross-modal reasoning questions (Cross-Modal QA), person understanding questions (Person QA), and questions about general knowledge (Knowledge QA).
@@ 3.1 M3-Bench-robot
Video Filming Recording videos with actual robots poses significant challenges due to high operational costs, hardware limitations, and deployment complexities. To address these constraints, we adopt a practical alternative: employing human actors to simulate robot behavior. This approach simplifies data collection while preserving both the first-person robot perspective and the multimodal quality required for our benchmark.
Each script involves multiple actors, with one designated to simulate the robot. This actor wears head-mounted camera equipment to capture the robot’s egocentric visual and auditory perspective. The resulting footage constitute the final videos in M3-Bench-robot. To ensure diversity and minimize location bias, we recruit 67 actors and film across 51 distinct locations, with no more than three videos recorded at each location.
We collect two types of audio tracks for each video. The first is directly recorded by the head-mounted device, reflecting the raw auditory input a robot would naturally receive, including ambient sounds and spatial acoustic variations. The second is captured using individual lapel microphones worn by each actor, providing high-fidelity voice recordings to complement the primary audio stream.
Annotations After recording the videos, annotators curate QA pairs for each video. Although some questions are pre-scripted, the final video content may deviate from the original script due to realistic filming conditions. Consequently, not all scripted questions remain applicable. Annotators carefully review each scripted question to determine whether it should be retained, revised, or discarded, and provide corresponding answers when necessary. For all retained or revised questions, annotators are required to specify the precise timestamp at which the question should be asked. Importantly, the timestamp must precede the robot’s corresponding response or action to avoid inadvertently revealing the answer.
In addition to the script-based questions, annotators are also required to create new questions to ensure that each video contains at least 12 QA pairs. All newly added questions should also align with one or more of the question types listed in Table 1.
Besides QA pair creation, annotators also generate subtitles to enhance the usability of the dataset. Specifically, they manually annotate the start and end timestamps for each dialogue segment, together with the speaker’s identity and the transcribed dialogue content.
@@ 3.2 M3-Bench-web
Video Collection The video collection adopts a question-driven approach: annotators select videos that could support the design of at least five questions belong to the types listed in Table 1. This strategy naturally leads to the selection of videos with rich narratives and complex inter-entity relationships, making them well-suited for assessing agent’s capability of reasoning with long-term memory.
To promote video diversity and avoid overrepresentation of easily annotated content, we provide annotators with a reference list of video categories emphasizing high information density and relevance to real-world multimodal agent applications. Annotators are required to submit up to 20 videos from each category and are allowed to suggest new categories, which are included if deemed sufficiently distinct from the existing category list by the authors. The final dataset comprises 46 distinct video types, as summarized in Figure 3.
QA Annotations The same annotator who collects the video also generates at least five corresponding question-answer pairs. Each question must correspond to at least one type defined in Table 1. In M3-Bench-web, all question timestamps are set to the end of the video. All questions are required to be specific, objective, and have a single unambiguous answer that can be reasonably derived from clues in the video, ensuring both the effectiveness and fairness of subsequent evaluation. For example, questions answerable from multiple perspectives or with ambiguous references, such as "the man" or "in the middle part of the video," are not considered valid. Appendix 9 provides the full annotation guidelines, annotators’ information, and quality control details for M3-Bench-web.
@@ 3.3 Automatic Evaluation
We use GPT-4o as an automatic evaluator for M3-Bench by prompting it to assess the correctness of a generated answer by comparing it to the corresponding reference answer for the same question. The prompt template is shown in Table LABEL:prompt_gpt4o_evaluation15.1).
@@ 4 Approach
search_clip Return top-kk most relevant memory clips for the given query. A memory clip refers to the agent’s episodic and semantic memory of a segment (typically 30 seconds) generated during clip-by-clip streaming video processing.
@@ 4.1 Long-Term Memory
Long-term memory is implemented as an external database that stores information in a structured, multimodal format (text, images, audio). Specifically, Memories are organized as an entity-centric multimodal graph, where each node represents a distinct memory item. Each node includes a unique ID, modality type, raw content, weight, embeddings, and other metadata such as timestamps. See Table 3 for details. Nodes are connected by undirected edges that represent logical relationships between memory items. For example, items sharing the same entity ID are linked to form an entity-centric memory graph. This design supports not only sequential retrieval of memories based on timestamps but also associative retrieval based on entities.
The agent constructs its memory by incrementally adding new text, image, or audio nodes. When a memory generated by the memorization process already exists in long-term memory, the corresponding node or edge is reactivated and its weight increased; if it is new, a corresponding node or edge is added to the graph. Conflicting information may be introduced during construction. To resolve this, M3-Agent applies a weight-based voting mechanism during inference: frequently activated entries accumulate higher weights and override conflicting entries with lower weights. This mechanism ensures the robustness and consistency of the memory graph over time.
Search Tool To facilitate memory retrieval, we provide a suite of search tools that enable the agent to retrieve relevant memories based on specific requirements. In particular, we implement two types of search mechanisms operating at different levels of granularity, as summarized in Table 4. Detailed implementation of these retrieval mechanisms is provided in Appendix 10.
@@ 4.4 Training
We apply reinforcement learning to optimize the M3-Agent. Although the memorization and control are conceptually handled by a single model, we trained two separate policy models to achieve optimal performance. Memorization relies strong multimodal understanding, while control requires strong reasoning capabilities. Accordingly, we initialized each policy model with different foundation models: Qwen2.5-Omni [51], an advanced open-source multimodal model supporting both visual and audio inputs, for memorization; and Qwen3 [53], an open-source large language model with powerful reasoning abilities, for control.
@@
Memorization To improve the model’s ability to generate desired memory, we perform imitation learning on Qwen2.5-Omni-7b to create memory-7b-sft. The process begins with constructing a high-quality synthetic demonstration dataset. We segment each video in the dataset into 30-second clips, and corresponding memory annotations are generated through a three-stage process: (1) Episodic memory synthesis: We perform a hybrid annotation strategy by jointly prompting Gemini-1.5-Pro and GPT-4o. Accordingly, GPT‑4o supplies frame‑level cues, which serve as priors for Gemini‑1.5‑Pro; the two outputs are merged to form richer narrative summaries than either alone. (2) Identity equivalence detection: We propose an algorithm that automatically mines high-confidence meta-clips, short monologue clips containing exactly one face and one voice, from a long video to construct a global face-voice correspondence. These meta-clips offer clear identity cues, enabling accurate face-voice pairing. Once the global mapping is established, it can be used to automatically annotate face-voice associations in any 30-second subclip. (3) Other semantic memory synthesis: We design prompt templates to extract semantic memories from various perspectives, guiding semantic memories to include information listed in Table 1011). Details of the data synthesis process are provided in Appendix 11. In total, we synthesize 10,952 samples: 10,752 for training and 200 for validation.
@@ 5.1 Baselines
We evaluate M3-Agent against three types of baselines:
@@
  • Gemini-Agent: Gemini-1.5-Pro is prompted separately for memorization and control process. During memorization, it receives the full video with audio, facial recognition results and speaker identification results to generate episodic and semantic memories, denoted as memory-gemini-prompt. In the control, it performs memory searches and generates responses, referred to as control-gemini-prompt.
  • Gemini-GPT4o-Hybrid: We also evaluate a setup where GPT-4o is prompted to perform memory search and generate responses (control-gpt4o-prompt). The memorization remains handled by memory-gemini-prompt.
@@ 5.3 Main Results
Table 5: Results on M3-Bench-robot, M3-Bench-web, and VideoMME-long. We also present a comparison of all methods across different question types in M3-Bench: multi-evidence reasoning (ME), multi-hop reasoning (MH), cross-modal reasoning (CM), person understanding (PU), and general knowledge extraction (GK).
@@
We also assess the memorization model via precision and comprehension, as reported in Appendix 12.
@@ 5.4 Ablation Study
To evaluate the impact of memorization on overall performance, we fixed the control model to control-7b-rl and compared different memorization methods, as shown in Table 6. First, we replaced the memory with that generated by memory-gemini-prompt, resulting in accuracy drops of 2.0%, 2.6%, and 9.1% on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. This suggests that memory-7b-sft produces higher-quality memory than memory-gemini-prompt. Next, we evaluated memory-7b-prompt, which led to accuracy reductions of 5.4%, 9.0%, and 11.0% on the same benchmarks, highlighting the importance of imitation learning in generating effective memory. Finally, we ablated key components in the memory generation process. The results show that removing character identity equivalence or semantic memory significantly degrades QA performance.
@@
Next, we investigate the impact of control on final performance. We fix memorization model as memory-7b-sft and evaluate various control models, as shown in Table 7. First, we compare two RL algorithms: GRPO and DAPO. Training details for GRPO are provided in Appendix 13. Our results show that control-32b-rl trained with DAPO consistently outperform control-32b-grpo across all test sets. Second, we analyze how DAPO’s performance scales with model size. The results indicate substantial improvements across all sizes. Specifically, after DAPO training, control-32b-rl achieves improvements of 10.0%, 8.0%, and 9.3% in accuracy over control-32b-prompt on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. Finally, we ablate two designs: inter-instruction and reasoning. Both are shown to be critical. Removing inter-instruction results in accuracy drops of 10.5%, 5.8%, and 5.9% on M3-Bench-robot, M3-Bench-web, and VideoMME-long, respectively. Removing reasoning leads to decreases of 11.7%, 8.8%, and 9.5% on the same benchmarks.